# Bayesian Optimization in Language space: An Eval-Efficient AI Self-Improvement Framework

### v1

@@ 7.1 Problem definition
Formally, define the brand fine-tuned ad-generation AI system as $\Phi$, whose behavior is controlled by a prompt set $\Pi$ (Figure 6). For a given prompt set $\Pi$, we write $\Phi(\Pi)$ for the corresponding ad creative shown on the ad platform. As a system, $\Phi(\Pi)$ maps the input space $\mathcal{X}$ (potential ad viewers on the platform) to a distribution over the output space $\mathcal{Y}$ (observable ad responses). For each $x\in\mathcal{X}$, we denote by $y(\Phi(\Pi),x)$ the random $\mathcal{Y}$-valued output of the system at input $x$. In a digital marketing setting, $x$ is a platform user, and a realization of $y(\Phi(\Pi),x)$ represents that user’s reaction (e.g., clicks, conversions, spend) to the ad $\Phi(\Pi)$. When a user $x\in\mathcal{X}$ is shown the ad creative $\Phi(\Pi)$, we observe an outcome $y\in\mathcal{Y}$, which is a realization of $y(\Phi(\Pi),x)$.
As discussed earlier, the distribution of exposed users depends on the ad $\Phi(\Pi)$ and therefore on the prompt set $\Pi$. We denote by $\mathcal{D}_{\mathcal{X}}(\Pi)$ the unknown distribution of users $x$ who are shown the ad generated from $\Pi$ when the ad $\Phi(\Pi)$ is run on the platform.
@@ 7.2 Application of T-BoN BO
$\pi_{t,g}^{j,(i)}\leftarrow\operatorname{Apply}\big(\pi_{t,g-1}^{j},\delta_{t,g}^{j,(i)}\big),\qquad i=1,\dots,N.$ (19)
@@
After all $J$ trajectories have been evaluated at iteration $t$, the algorithm updates the global meta-reflection $R_{t}=\textsc{MetaReflect}(H_{t})$, as in Section 5.3. Concretely, $M_{\text{critic}}$ is prompted to review a subset of the best- and worst-performing ads and prompts in $H_{t}$ and to summarize the key patterns and rules that distinguish successful creatives from unsuccessful ones (see Figure 13 in Appendix A). For example, the reflection may emphasize that scenes with social eating and visible grill marks perform better than isolated product shots, or that emphasizing “satisfying” and “juicy” yields higher effectiveness scores. This reflection is then injected into the context of subsequent textual-gradient generation and Best-of-$N$ comparisons, allowing all trajectories to share accumulated knowledge even though they explore different regions of the prompt space.
@@ 8 Experiments
We validate the empirical effectiveness of the proposed T-BoN BO framework for a real-world digital marketing problem described in 7. Specifically, we examine if and how much T-BoN BO improves outcomes under a fixed evaluation budget compared to state-of-the-art baselines.
@@ 8.2 Experiment Setup
For each scenario, from its creative brief, we generated 64 high-quality initial prompts that were used as input to the state-of-the-art image generation AI model named Imagen 4  [^12] See Vertex AI’s description (https://console.cloud.google.com/vertex-ai/publishers/google/model-garden/imagen-4.0-generate-preview-06-06) for details. to generate the initial 64 high-quality ads. (For prompts, see Figure 14 and 15 in Appendix A.) To maintain a rigorous baseline, we made the 64 initial prompts detailed enough to yield strong ads. This ensures that the algorithm’s subsequent improvements are not explained as trivial quality gains from having initial under-specified prompts. That is, the measured improvements are attributed to the algorithm’s ability to align ads with the underlying persona distribution better.
@@ 8.5 Ablation study
The main experiment in Section 8.1-8.4 evaluates the candidate ad at each iteration as the mean of scores across 411 personas (20% of the persona data in Twin-2k-500 (toubia2025twin)) that were injected into the LLM’s context. Each persona consists of a large amount of information, with 500 survey questions accounting for around 50,00050{,}000 tokens. Would T-BoN BO still outperform strong state-of-the-art baselines when the LLM receives less persona context information?

### v2

@@ 4.3 Theory: Best-of-N gradient’s Asymptotic Equivalence to UCB gradient
5 TextGrad-Best-of-N Bayesian Optimization (T-BoN BO)
We now describe the self-improving AI framework proposed in this paper: TextGrad-Best-of-N Bayesian Optimization (T-BoN BO). As discussed in Section 1, T-BoN BO is designed to implement parallel gradient-based UCB-BO in the discrete space of natural language (Figure 3). Specifically, T-BoN BO implements the parallel gradient-based UCB-BO that reuses the final gradient point as the initialization for the next iteration. Therefore, each initial point gives rise to a trajectory: a sequence of candidate points obtained by repeatedly taking gradient-ascent steps and evaluations starting from that initialization.
[…]
Figure 4: Visual summary of T-BoN BO (Algorithm 1) with $T$ iterations, $J$ trajectories, and $G$ Best-of-NN gradients per each iteration and trajectory.
Algorithm.
Throughout $T$ number of optimization iterations, T-BoN BO runs $J$ trajectories in parallel, starting from a set of initial prompts, denoted $\{\Pi_{0}^{j}\}_{j=1}^{J}$. At each iteration and a trajectory, we fix the hyperparameters: the number of candidates per gradient step ($N$) and the number of gradient steps per iteration ($G$).
At iteration $t=0$, it starts from evaluating $\{\Pi_{0}^{j}\}_{j=1}^{J}$, i.e., the initial systems $\{c_{0}^{j}=\Phi(\Pi_{0}^{j})\}_{j=1}^{J}$, on $\{\mathcal{D}_{\mathcal{X}}(\Pi_{0}^{j})\}_{j=1}^{J}$ to obtain initial scores $\{s_{0}^{j}\}_{j=1}^{J}$. The history is then initialized as $H_{0}=\bigcup_{j=1}^{J}\{(\Pi_{0}^{j},c_{0}^{j},s_{0}^{j})\}$. The shared reflection is initialized as $R_{0}=\emptyset$.
At each iteration $t$, for each trajectory $j\in\{1,\ldots,J\}$, we set $\Pi_{t,0}^{j}$ as $\Pi_{t-1}^{j}$ and proceed as follows:
  • 1.
    Best-of-N gradients. (Detailed in Section 4.2) For each trajectory $j$, repeat steps 1) – 4) for $g=1,\ldots,G$:
    1. 1)
      Textual gradient generation: From the current prompt $\Pi_{t,g-1}^{j}$, query the textual-gradient generation operator $\nabla_{\text{text}}(\Pi_{t,g-1}^{j};R_{t-1},M_{\text{critic}})$ independently $N$ times to obtain textual gradients $\{\delta^{j,(i)}_{t,g}\}_{i=1}^{N}$.
    2. 2)
      Candidate prompt generation: Apply each textual gradient $\delta_{t,g}^{j,(i)}$ to $\Pi_{t,g-1}^{j}$ for all $i$ to get $\{\Pi_{t,g}^{j,(i)}\}_{i=1}^{N}$.
    3. 3)
      Candidate system outcome generation: Produce the system outcomes $\{c_{t,g}^{j,(i)}:=\Phi(\Pi_{t,g}^{j,(i)})\}_{i=1}^{N}$.
    4. 4)
      Best-of-N selection: Find the best of $\{(\Pi_{t,g}^{j,(i)},c_{t,g}^{j,(i)})\}_{i=1}^{N}$ using the critic and the shared reflection $R_{t-1}$, and set the winner as $(\Pi_{t,g}^{j},c_{t,g}^{j})$.
  • 2.
    Evaluation. For each trajectory $j$,
    1. 1)
      Set $c_{t}^{j}=c_{t,G}^{j}$ and $\Pi_{t}^{j}=\Pi_{t,G}^{j}$.
    2. 2)
      Evaluate $c_{t}^{j}$ to get the score $s_{t}^{j}=\mathrm{Eval}(c_{t}^{j},\mathcal{D}_{\mathcal{X}}(\Pi_{t}^{j}))$, where
      $\mathrm{Eval}(c_{t}^{j},\mathcal{D}_{\mathcal{X}}(\Pi_{t}^{j}))\;:=\;\frac{1}{L}\sum_{l=1}^{L}r\!\left(y\left(c^{j}_{t},x_{t}^{j,(l)}\right)\right),\qquad x_{t}^{j,(l)}\sim\mathcal{D}_{\mathcal{X}}(\Pi^{j}_{t}),$ (22)
    3. 3)
      Given $\{(\Pi_{t}^{j},c_{t}^{j},s_{t}^{j})\}_{j=1}^{J}$, append it to the shared history $H_{t-1}$ to form $H_{t}$.
    4. 4)
      If $s_{t}^{j}>s_{t-1}^{j}$, keep $(\Pi_{t}^{j},c_{t}^{j})$. Otherwise, revert trajectory $j$ to the previous state: $(\Pi_{t}^{j},c_{t}^{j})=(\Pi_{t-1}^{j},c_{t-1}^{j})$ and set $s_{t}^{j}=s_{t-1}^{j}$.
    For 4), we follow textgrad and adopt a simple acceptance rule that controls progress: if $s_{t}\geq s_{t-1}$, the new candidate is adopted; otherwise, the system rolls back to the previous iteration’s outcome, i.e., $(\Pi_{t-1},c_{t-1},s_{t-1})$. After evaluation of all trajectories, $\{(\Pi^{j}_{t},c^{j}_{t},s^{j}_{t})\}_{j=1}^{J}$ are appended to the history $H_{t-1}$ to form $H_{t}$.
  • 3.
    Shared Meta-reflection.
    At the end of each iteration $t$, we update the global reflection
    $R_{t}\leftarrow\textsc{MetaReflect}(H_{t};M_{\text{critic}})$
    using the aggregated history from all trajectories for use at iteration t+1t{+}1. Here, $\textsc{MetaReflect}(H_{t};M_{\text{critic}})$ compresses the full history $H_{t}$ of prompts, generated system outcomes, and evaluation results into a natural language reflection $R_{t}$. $R_{t}$ consists of a small set of actionable rules. In the ad generation example, those rules may take the form “emphasize social context,” “avoid cluttered backgrounds,” or “prefer warm lighting.” This reflection serves as a global memory; in subsequent iterations, the critic LLM takes this reflection as a part of its context when generating textual gradients. (For the prompt used for meta-reflection, see Figure 12 in Appendix A).
After $T$ optimization steps, the sequence of best-performing outcomes $\{c_{t}^{j_{t}^{\ast}}\}_{t=1}^{T}$, where $j_{t}^{\ast}:=\operatorname{argmax}_{j}s_{t}^{j}$, is reported as the algorithm’s result.
Note that T-BoN BO takes $N\times G$ generation of prompt/outcome candidates before each evaluation. This careful selection of a prompt/outcome candidate to evaluate is the key to evaluation efficiency in T-BoN BO: it deliberately trades inexpensive generation compute for fewer evaluations.
Input: System $\Phi$; $\{\Pi_{0}^{j}\}_{j=1}^{J}$; critic $M_{\text{critic}}$; candidates per step $N$; gradient steps $G$; iterations $T$; $J$ trajectories Output: Optimized triples $\{(\Pi^{j}_{T},c^{j}_{T},s^{j}_{T})\}_{j=1}^{J}$ 1 for $j=1$ to $J$ do 2 $c_{0}^{j}\leftarrow\Phi(\Pi_{0}^{j})$; 3 $s_{0}^{j}\leftarrow\text{Eval}(c_{0}^{j},\mathcal{D}_{\mathcal{X}}(\Pi_{0}))$; 4 $H_{0}\leftarrow\{(\Pi_{0}^{j},c_{0}^{j},s_{0}^{j})\}$, $R\leftarrow\emptyset$; 5  6 end for 7 8for $t=1$ to $T$ do 9 for $j=1$ to $J$ do 10    $\Pi^{j}_{t,0}\leftarrow\Pi^{j}_{t-1}$; 11    for $g=1$ to $G$ do 12       for $i=1$ to $N$ do 13          $\delta_{t,g}^{j,(i)}\leftarrow\nabla_{\text{text}}(\Pi_{t,g-1}^{j};R,M_{\text{critic}})$ ; 14          $\Pi_{t,g}^{j,(i)}\leftarrow\operatorname{Apply}(\Pi_{t,g-1}^{j},\delta_{t,g}^{j,(i)})$ ; 15          $c_{t,g}^{j,(i)}\leftarrow\Phi(\Pi_{t,g}^{j,(i)})$ ; 16           17         end for 18       $(\Pi_{t,g}^{j},c_{t,g}^{j})\leftarrow\textsc{Best-of-N}(\{(\Pi_{t,g}^{j,(i)},c_{t,g}^{j,(i)})\},R_{t-1},M_{\text{critic}})$; 19        20      end for 21    $c_{t}^{j}\leftarrow c^{j}_{t,G}$, $\Pi_{t}^{j}\leftarrow\Pi_{t,G}^{j}$; 22    $s_{t}^{j}\leftarrow\text{Eval}(c_{t}^{j},\mathcal{D}_{\mathcal{X}}(\Pi_{t}))$; 23    $H_{t}\leftarrow H_{t-1}\cup\{(\Pi_{t}^{j},c_{t}^{j},s_{t}^{j})\}$; 24    if $s_{t}^{j}<s_{t-1}^{j}$ then 25       $\Pi_{t}^{j}\leftarrow\Pi_{t-1}^{j}$, $c_{t}^{j}\leftarrow c_{t-1}^{j}$, $s_{t}^{j}\leftarrow s_{t-1}^{j}$; 26        27      end if 28     29   end for 30 $R_{t}\leftarrow\textsc{MetaReflect}(H_{t})$; 31  32 end for return $\{c_{t}^{j_{t}^{*}}\}_{t=1}^{T}$, where $j_{t}^{*}:=\operatorname{argmax}_{j}s_{t}^{j}$ Algorithm 1 T-BoN BO
@@ 6.1 Problem definition
Formally, define the brand fine-tuned ad-generation AI system as $\Phi$, whose behavior is controlled by a prompt $\Pi$ (Figure 5). For a given prompt $\Pi$, we write $\Phi(\Pi)$ for the corresponding ad creative shown on the ad platform. As a system, $\Phi(\Pi)$ maps the input space $\mathcal{X}$ (potential ad viewers on the platform) to a distribution over the output space $\mathcal{Y}$ (observable ad responses). For each $x\in\mathcal{X}$, we denote by $y(\Phi(\Pi),x)$ the random $\mathcal{Y}$-valued output of the system at input $x$. In a digital marketing setting, $x$ is a platform user, and a realization of $y(\Phi(\Pi),x)$ represents that user’s reaction (e.g., clicks, conversions, spend) to the ad $\Phi(\Pi)$. When a user $x\in\mathcal{X}$ is shown the ad creative $\Phi(\Pi)$, we observe an outcome $y\in\mathcal{Y}$, which is a realization of $y(\Phi(\Pi),x)$.
As discussed earlier, the distribution of exposed users depends on the ad $\Phi(\Pi)$ and therefore on the prompt $\Pi$. We denote by $\mathcal{D}_{\mathcal{X}}(\Pi)$ the unknown distribution of users $x$ who are shown the ad generated from $\Pi$ when the ad $\Phi(\Pi)$ is run on the platform.
@@ 6.2 Application of T-BoN BO
$\pi_{t,g}^{j,(i)}\leftarrow\operatorname{Apply}\big(\pi_{t,g-1}^{j},\delta_{t,g}^{j,(i)};M_{\text{critic}}\big),\qquad i=1,\dots,N.$ (26)
@@
After all $J$ trajectories have been evaluated at iteration $t$, the algorithm updates the global meta-reflection $R_{t}=\textsc{MetaReflect}(H_{t})$, as in Section 5. Concretely, $M_{\text{critic}}$ is prompted to review a subset of the best- and worst-performing ads and prompts in $H_{t}$ and to summarize the key patterns and rules that distinguish successful creatives from unsuccessful ones (see Figure 12 in Appendix A). For example, the reflection may emphasize that scenes with social eating and visible grill marks perform better than isolated product shots, or that emphasizing “satisfying” and “juicy” yields higher effectiveness scores. This reflection is then injected into the context of subsequent textual-gradient generation and Best-of-$N$ comparisons, allowing all trajectories to share accumulated knowledge even though they explore different regions of the prompt space.
@@ 7 Experiments
We validate the empirical effectiveness of the proposed T-BoN BO framework for a real-world digital marketing problem described in 6. Specifically, we examine if and how much T-BoN BO improves outcomes under a fixed evaluation budget compared to state-of-the-art baselines.
@@ 7.2 Experiment Setup
For each scenario, from its creative brief, we generated 64 high-quality initial prompts that were used as input to the state-of-the-art image generation AI model named Imagen 4  [^8] See Vertex AI’s description (https://console.cloud.google.com/vertex-ai/publishers/google/model-garden/imagen-4.0-generate-preview-06-06) for details. to generate the initial 64 high-quality ads. (For prompts, see Figure 13 and 14 in Appendix A.) To maintain a rigorous baseline, we made the 64 initial prompts detailed enough to yield strong ads. This ensures that the algorithm’s subsequent improvements are not explained as trivial quality gains from having initial under-specified prompts. That is, the measured improvements are attributed to the algorithm’s ability to align ads with the underlying persona distribution better.
@@ 7.5 Ablation study
The main experiment in Section 7.1-7.4 evaluates the candidate ad at each iteration as the mean of scores across 411 personas (20% of the persona data in Twin-2k-500 (toubia2025twin)) that were injected into the LLM’s context. Each persona consists of a large amount of information, with 500 survey questions accounting for around 50,00050{,}000 tokens. Would T-BoN BO still outperform strong state-of-the-art baselines when the LLM receives less persona context information?